How many moles of carbon dioxide gas is produced when 20.5 grams of ethyl alcohol is burned?


1
Expert's answer
2020-12-03T14:01:36-0500

Solution:

The balanced chemical equation is:

C2H5OH(l) + 3O2(g) = 2CO2(g) + 3H2O(l)


According to the equation: Moles of C2H5OH = Moles of CO2 / 2


Moles of C2H5OH = Mass of C2H5OH / Molar mass of C2H5OH

The molar mass of C2H5OH is 46.07 g mol-1.

Hence,

Moles of C2H5OH = 20.5 g / 46.07 g mol-1 = 0.445 mol


Moles of CO2 = 2 × Moles of C2H5OH = 2 × 0.445 mol = 0.89 mol

0.89 moles of CO2


Answer: 0.89 moles of carbon dioxide (CO2) gas is produced.